# Is attention all you need to solve the correlated electron problem?

### v2

@@ Abstract
The attention mechanism has transformed artificial intelligence research by its ability to learn relations between objects. In this work, we explore how a many-body wavefunction ansatz constructed from a large-parameter self-attention neural network can be used to solve the interacting electron problem in solids. By a systematic neural-network variational Monte Carlo study on a moiré quantum material, we demonstrate that the self-attention ansatz provides an accurate, efficient, and unbiased solution. Moreover, our numerical study finds that the required number of variational parameters scales roughly as $N^{2}$ with the number of electrons, which opens a path towards efficient large-scale simulations.
@@ I Introduction
In the remainder of the article, we detail the self-attention neural network wavefunction ansatz and present its benchmark with self-consistent Hartree-Fock (scHF) and band-projected exact diagonalization. First, in Sec. II we summarize the traditional methods of Hartree-Fock and band-projected exact diagonalization for comparison. Then, Sec. III describes the construction of the self-attention neural network ansatz. Sec. IV summarizes principal ideas of the variational Monte Carlo method applied to optimize the wavefunction ansatz. Finally, our numerical results and benchmarks are presented in Sec. V.
@@ III Neural Network Wavefunction
Figure 1: Architecture of the neural network wavefunction ansatz. SlaterNet: Multilayer perceptron neural network generates one-body orbitals to approximate general single Slater determinant wavefunctions. Self-attention NN: Self-attention neural network for solids based on Psiformer [11] capturing correlations by mixing individual particle streams.
@@
In the following, we show how an expressive neural network wavefunction ansatz capturing correlations between particles can be constructed. First, in Sec. III.1, a universal approximation of single-particle orbitals is constructed by a deep feed-forward neural network. Second, Sec. III.2, by generating a set of single-particle orbitals from the feed-forward neural network, a single Slater determinant wavefunction is obtained, which enables neural network implementation of unrestricted Hartree Fock calculations. Finally, Sec. III.3, electron correlations are captured by mixing different electron streams in a permutation-equivariant way with self-attention mechanism. The resulting neural network wavefunction takes the form of generalized Slater determinants that are built from “correlated orbitals”, and by construction, is antisymmetric in electron coordinates [10]. Fig. 1 shows schematically and compares the network architectures representing general single-Slater determinant wavefunctions, and its extension by self-attention capturing correlations.
@@ III.1 Deep neural network for one-body orbitals
for any particle $i$, where the two vectors $\bm{L}=n\bm{L}_{1}+m\bm{L}_{2}$ with $n,m\in\mathbb{Z}$ specify the supercell size and geometry. For periodic systems, we choose the function ${\rm feature}(\bm{r})$ to express the coordinate $\bm{r}$ in terms of periodic coordinates [pescia2022neural, 40],
@@ III.2 SlaterNet: Neural network for unrestricted Hartree-Fock
where the same transformations $W^{l}$ and shifts $\bm{b}^{l}$ are applied to the stream generated by the coordinates of each particle. Thereby, this compositional function generates a single function in a high-dimensional space before projecting onto the space of $N$ single-particle orbitals $\mathbb{C}^{N}$. Then, $N$ distinct single-particle orbitals are constructed as
@@ IV.2 Optimizing variational parameters
To optimize the parameters of the wavefunction, the steepest descent direction is identified as the direction that minimizes the cost function $L[\bm{\theta}+d\bm{\theta}]$ for a step $d\bm{\theta}$, where the distance of the step is measured on the variational wavefunction space $||d\bm{\theta}||_{\Psi}$ [49],
@@ V.1 Convergence and scaling with system size
$N_{\text{par}}^{*}\approx 400\times N^{2.0}\ .$ (32)
The fit is shown in Fig. 4. This empirical result suggests a significantly better scaling compared to tensor network methods whose required number of parameters scales as $\propto e^{\sqrt{N}}$ [53].
While this scaling law provides a useful estimate of the number of parameters required to achieve a given accuracy, the mere parameter count is not the only relevant factor. Since attention heads play a crucial role in capturing correlations, a minimum number of attention heads and layers (approximately above 3 in our tests) is necessary to reach the ground state. Simply increasing the total number of parameters without ensuring sufficient attention heads and layers would not improve performance. Our scaling law serves as a guideline within the regime where these minimal requirements are met. We leave a more detailed analysis of scaling laws as a function of individual network dimensions for future work.
Figure 4: Scaling law for the number of variational parameters at convergence saturation as a function of the number of electrons for $\epsilon=10$. The extracted saturation points are well described by a power-law relation indicating how the required parameter count grows with system size. This scaling provides a guideline for selecting an optimal network size while ensuring convergence.
@@ V.2 Benchmark for small system size
After ensuring convergence and saturation in the number of variational parameters, we benchmark our self-attention NN results with scHF and BP-ED to assess the quantitative accuracy of predictions from the self-attention NN.
We study a variety of different system sizes and regimes at different interaction strengths. The comprehensive energy comparisons given in Fig. 5 graphically evince the difference in ground state energy levels between all 3 methods for 9 unit cell system again at $\nu=2/3$. The self-attention NN energy is considerably lower than scHF, with the difference precisely representing the correlation energy. We see that the correlation energy is about $2\%$ in our moiré system, considerably higher than the correlation energy in ordinary molecular systems [3]. This enhanced correlation energy underscores the relevance of moiré systems for studying strongly correlated phenomena. Additionally, we observe that as the number of bands increases in BP-ED, the ground-state energy approaches but remains higher than that of the self-attention NN .
The advantage of self-attention NN becomes even more pronounced in larger systems, such as those with 18 electrons [Table 1], where BP-ED is limited to a single band due to the enormous Hilbert space dimension. In this case, the self-attention NN energy is nearly $2.5\%$ lower than that of BP-ED. In fact, the band mixing is so strong that even scHF achieves a lower energy than BP-ED.
These results demonstrate that self-attention NN consistently achieves lower energies than both BP-ED and scHF, underscoring its capacity to express ground states of the moiré systems to higher accuracy, and revealing the variational superiority of our self-attention NN method.
@@ V.3 Fermi Liquids and Generalized Wigner Crystals
Finally, the difference in energies from the self-attention NN and scHF allows to estimate correlation energy Eq. (3). We include the extracted correlation energy as a function of dielectric constant in App. D. Furthermore, in App. E, we provide additional exact-diagonalization results on the metal-insulator phase transition.
@@ VII Conclusion
The conceptual strength of the self-attention neural network ansatz lies in its two pillars of construction: (i) it employs a large scale parameterization without human bias, (ii) the self-attention mechanism efficiently learns relations between electrons that encode their correlations. Altogether, these results put faith in the hope that a unified wavefunction ansatz can accurately describe a wide range of, if not all, quantum phases in strongly interacting electron systems.
@@ Appendix A Coulomb system with periodic boundary condition
$\ +\frac{2\pi}{A_{{\rm u.c.}}}\sum_{\bm{G}\neq 0}e^{i\bm{G}(\bm{r}-\bm{r}_{a})}\frac{{\rm Erfc}\left(\eta G\right)}{G}-\xi_{0}^{{\rm L}}$ (44)

### v3

@@ Abstract
The attention mechanism has transformed artificial intelligence research by its ability to learn relations between objects. In this work, we explore how a many-body wavefunction ansatz constructed from a large-parameter self-attention neural network can be used to solve the interacting electron problem in solids. By a systematic neural-network variational Monte Carlo study on a moiré quantum material, we demonstrate that the self-attention ansatz provides an accurate and efficient solution without human bias. Moreover, our numerical study finds that the required number of variational parameters scales roughly as $N^{2}$ with the number of electrons, which opens a path towards efficient large-scale simulations.
@@ I Introduction
In the remainder of the article, we detail the self-attention neural network wavefunction ansatz and present its benchmark with self-consistent Hartree-Fock and band-projected exact diagonalization. First, in Sec. II we summarize the traditional methods of Hartree-Fock and band-projected exact diagonalization for comparison. Then, Sec. III describes the construction of the self-attention neural network ansatz. Sec. IV summarizes principal ideas of the variational Monte Carlo method applied to optimize the wavefunction ansatz. Finally, our numerical results and benchmarks are presented in Sec. V.
@@ III Neural Network Wavefunction
Figure 1: Architecture of the neural network wavefunction ansatz. SlaterNet: Multilayer perceptron neural network generates one-body orbitals to approximate general single Slater determinant wavefunctions. Self-attention NN: Self-attention neural network for solids based on Psiformer [11] capturing correlations by communicating mutual information between individual particle streams.
@@
In the following, we show how an expressive neural network wavefunction ansatz capturing correlations between particles can be constructed. First, in Sec. III.1, a universal approximation of single-particle orbitals is constructed by a deep feed-forward neural network. Second, Sec. III.2, by generating a set of single-particle orbitals from the feed-forward neural network, a single Slater determinant wavefunction is obtained, which enables neural network implementation of unrestricted Hartree Fock calculations. Finally, Sec. III.3, electron correlations are captured by mixing different electron streams in a permutation-equivariant way with self-attention mechanism. The resulting neural network wavefunction takes the form of generalized Slater determinants that are built from “correlated orbitals”, and by construction, is antisymmetric in electron coordinates [10]. This representation is in principle universal as all many-body fermionic wavefunctions can be written as a generalized Slater determinant when allowing for discontinuous functions [10]. Fig. 1 shows schematically and compares the network architectures representing general single-Slater determinant wavefunctions, and its extension by self-attention capturing correlations.
@@ III.1 Deep neural network for one-body orbitals
for any particle $i$, where the two vectors $\bm{L}=n\bm{L}_{1}+m\bm{L}_{2}$ with $n,m\in\mathbb{Z}$ specify the supercell size and geometry. For periodic systems, we choose the function ${\rm feature}(\bm{r})$ to express the coordinate $\bm{r}$ in terms of periodic coordinates [40, 41],
@@ III.2 SlaterNet: Neural network for unrestricted Hartree-Fock
where the same transformations $W^{l}$ and shifts $\bm{b}^{l}$ are applied to the single-particle stream $\bm{h}_{i}^{l}$ generated by the coordinates of each particle. Thereby, this compositional function generates a single function in a high-dimensional space before projecting onto the space of $N$ single-particle orbitals $\mathbb{C}^{N}$. Then, $N$ distinct single-particle orbitals are constructed as
@@ IV.2 Optimizing variational parameters
To optimize the parameters of the wavefunction, the steepest descent direction is identified as the direction that minimizes the cost function $L[\bm{\theta}+d\bm{\theta}]$ for a step $d\bm{\theta}$, where the distance of the step is measured on the variational wavefunction space $||d\bm{\theta}||_{\Psi}$ [50, 51],
@@
We remark that around a second-order quantum phase transition, the Fisher information diverges in the parameter that drives the quantum phase transition [53, 54]. These divergencies may pose difficulties in the numerical approximation of the Fisher information matrix so that optimization around a quantum critical point may be more challenging than deep within a phase – even if the variational ansatz remains sufficiently expressive to describe the critical state.
@@ V.1 Convergence and scaling with system size
$N_{\text{par}}^{*}\approx(400\pm 49)\times N^{2.01\pm 0.05}\ .$ (32)
The fit is shown in Fig. 4. While this scaling behavior is derived for the present system and network architecture, we believe it may serve as a useful approximation for estimating parameter growth in other settings as well, though it may not directly generalize.
This scaling law provides a helpful estimate for the number of parameters required to achieve a given accuracy; however, the parameter count alone is not the only relevant factor. Since attention heads play a crucial role in capturing correlations, a minimum number of attention heads and layers (approximately above 3 in our tests) is necessary to reach the ground state. Simply increasing the total number of parameters without ensuring sufficient attention heads and layers would not improve performance. Our scaling law serves as a guideline within the regime where these minimal requirements are met. We leave a more detailed analysis of scaling laws as a function of individual network dimensions for future work.
In comparison, tensor network methods for two-dimensional lattices under area-law assumptions require a number of parameters scaling as $\propto e^{\sqrt{N}}$ [57]. For tensor networks, parameter scaling depends on the entanglement structure of the problem and tensor network ansatz. Our empirical result of quadratic scaling of the parameters in the self-attention NN ansatz suggests significantly better scaling compared to tensor network methods.
Figure 4: Scaling law for the number of variational parameters at convergence saturation as a function of the number of electrons for $\epsilon=10$. The extracted saturation points are well described by a power-law relation $N^{*}=C\cdot N^{\alpha}$, with best-fit parameters $\alpha=2.01\pm 0.05$ and $C=(4.00\pm 0.49)\times 10^{2}$. This relation quantifies how the required parameter count grows with system size and provides a practical guideline for selecting an optimal network size while ensuring convergence.
@@ V.2 Benchmark for small system size
After ensuring convergence and saturation in the number of variational parameters, we benchmark our self-attention NN results with SlaterNet and BP-ED to assess the quantitative accuracy of predictions from the self-attention NN.
We study a variety of different system sizes and regimes at different interaction strengths. The comprehensive energy comparisons given in Fig. 5 graphically evince the difference in ground state energy levels between all 3 methods for 9 unit cell system again at $\nu=2/3$. The self-attention NN energy is considerably lower than SlaterNet, with the difference precisely representing the correlation energy. We see that the correlation energy is about $2\%$ in our moiré system, considerably higher than the correlation energy in ordinary molecular systems [3]. This enhanced correlation energy underscores the relevance of moiré systems for studying strongly correlated phenomena. Additionally, we observe that as the number of bands increases in BP-ED, the ground-state energy approaches but remains higher than that of the self-attention NN .
The advantage of self-attention NN becomes even more pronounced in larger systems, such as those with 18 electrons [Table 1], where BP-ED is limited to a single band due to the enormous Hilbert space dimension. In this case, the self-attention NN energy is nearly $2.5\%$ lower than that of BP-ED. In fact, the band mixing is so strong that even SlaterNet achieves a lower energy than BP-ED.
These results demonstrate that self-attention NN consistently achieves lower energies than both BP-ED and SlaterNet, underscoring its capacity to express ground states of the moiré systems to higher accuracy, and revealing the variational superiority of our self-attention NN method.
@@ V.3 Fermi Liquids and Generalized Wigner Crystals
Finally, the difference in energies from the self-attention NN and SlaterNet allows to estimate correlation energy Eq. (3). We include the extracted correlation energy as a function of dielectric constant in App. D. Furthermore, in App. E, we provide additional exact-diagonalization results on the metal-insulator phase transition.
@@ VI Comparison to other neural network ansätze
We highlight that our results have been computed from a randomly initialized wavefunction. The only physical input is the determinant operation guaranteeing the antisymmetric structure of the fermionic wavefunction. In contrast, previous works [10, 11] applied pretraining with Hartree-Fock to initialize the wavefunction before optimizing energy. Also, recent message-passing architectures [23] describe correlations starting from a chosen Slater determinant of single-particle orbitals.
In the future, our work can be extended to include computation of excited states [60] and zero-temperature Green’s functions and spectral functions [61]. Finite temperature effects can be studied by an extension of the neural network architecture to represent thermal density matrices [62, 63]. Dynamical and open systems can be studied by applying the time-dependent variational principle [64] to a variational density matrix [65, 66, 67, 68, 69, 70].
@@ VII Conclusion
The conceptual strength of the self-attention neural network ansatz lies in its two pillars of construction: (i) it employs a large scale parameterization without human bias, (ii) the self-attention mechanism efficiently learns relations between electrons that encode their correlations. Altogether, these results put faith in the hope that a unified wavefunction ansatz may be able to accurately describe a wide range of, if not all, quantum phases in strongly interacting electron systems.
@@ Appendix A Coulomb system with periodic boundary condition
$\ +\frac{2\pi}{A_{{\rm u.c.}}}\sum_{\bm{G}\neq 0}\frac{{\rm Erfc}\left(\eta G\right)}{G}-\xi_{0}^{{\rm L}}$ (44)
@@ Appendix C Hyperparameters
For the self-attention neural network as described in Sec. III.3, the total number of variational parameters depends on the network dimensions as
$N_{\rm par}$ $=2d_{\rm dim}d_{\rm L}$
$+L\bigg{(}2N_{\rm heads}d_{\rm Attn}d_{\rm L}$
$\qquad+2N_{\rm heads}d_{\rm AttnValues}d_{\rm L}$
$\qquad+d_{\rm L}(d_{\rm L}+1)\bigg{)}$
$+2d_{\rm L}N_{\rm el}N_{\rm det}$
where the first line is the embedding $W_{0}$ of the periodic features in the layer dimension $d_{L}$, the second line are key and query matrices $W_{\rm k}^{lh},\ W_{\rm q}^{lh}$ for each layer $l=1,...,L$ and head $h=1,...,N_{\rm heads}$, the third line are the value matrices $W_{\rm k}^{lh}$ and projection matrices $W_{\rm o}^{l}$ onto the layer dimension, the fourth line are the perceptron weights and biases $W^{l},\bm{b}_{l}$, and the last line are the projection $\bm{w}_{2j}^{m},\bm{w}_{2j+1}^{m}$ on $j=1,...,N_{\rm el}$ complex orbitals for $m=1,...,N_{\rm det}$ determinants.